OpenStudy (anonymous):

f(x) = 2.5x – 10.5 in this equation, i subtract 2.5 from 10.5, correct? medal and fan! :)

OpenStudy (triciaal):

where is the rest of the question?

OpenStudy (anonymous):

well, it's part of a graph.

OpenStudy (triciaal):

all you have is the equation. It does not say what to do.

OpenStudy (anonymous):

i just need to know if i am to subtract 2.5 from 10.

OpenStudy (anonymous):

i have to complete a table to solve the equation 2.5x - 10.5 = 64(0.5^x)

OpenStudy (triciaal):

y = f(x) slope = 2.5 and b = -10.5 y = mx + b

OpenStudy (anonymous):

I don't know if that helps me.

OpenStudy (anonymous):

need to see the table? so you can see the full problem?

OpenStudy (triciaal):

from the table what value of x will make the equation true?

OpenStudy (anonymous):